# The discrete generalized exchange-driven system

### v1

@@ Abstract.
Corresponding author: F.P. da Costa
@@ 1. Introduction
In the present paper our primary goal is to propose a new, more general, discrete model for exchange-driven growth of cluster and start its mathematical study. This new model will be refered to as the discrete generalized exchange-driven growth model (DGED, for short). As explained in the next session, the DGED model allows for the exchange not only of a monomer between two interacting clusters, but of the transfer of a bigger chunk of one cluster to the other. Thus, the DGED model includes as a special case the EDG model referred to above. In a certain sense, the usual coagulation-fragmentation equations (with binary fragmentation) can also be considered as a special case of the DGED equations.
@@
The outline of this article is as follows. In sections 2 and 3, we introduce the DGED model for two cases that, in section 2, are called isolated and non-isolated, and some relevant mathematical notions and results that will be needed later in this work. In section 4 we consider a truncated finite dimensional version of the DGED system. This is a finite dimensional ordinary differential equation system, for which the existence, uniqueness, and non-negativity of solutions to Cauchy problems are automatically obtained from the standard theory. Here we prove some properties of its solutions that will play an important role in the next section. In section 5 we establish the existence of mild solutions to Cauchy problems for the DGED system, i.e., continuous solutions of the integral version of the differential equations system. This is done by the use of Helly’s theorem and a diagonal argument to establish the existence of a function which is the limit of a (sub)sequence of solutions to truncated systems as the dimension of the truncation grows to infinity, and then by proving that this function is indeed a (mild) solution of the DGED system using appropriate bounds (obtained in the previous two sections) and the bounded and monotone convergence theorems. Still in section 5 we prove that for the isolated case solutions obtained in this way (which are called admissible) conserve two important quantities, which have the “physical” interpretation of being the total number of clusters, and the total mass of the system, whereas in the non-isolated case only mass is conserved. In the short section 6 we prove that under slightly more restrictive conditions on the rate coefficients of the DGED system and assuming conservation of the total cluster mass (which is true for admissible solutions in both the isolated and non-isolated cases), then each component of the mild solution proved to exist in section 5 is indeed continuously differentiable. In Section 7, we state and prove a uniqueness result under a more restrictive growth condition on the rate coefficients, applying standard techniques used in coagulation-type systems. We conclude the paper with a section of final remarks where we discuss some possible future work related to the long-term behavior of solutions for this model.
@@ 2. The generalized exchange-driven growth model
Figure 1. Reaction scheme of the DGED model considered in this paper. A chunk of size $k$ in a particle of size $p$ is transferred to a particle of size $q$ to produce a particle of size $p-k$ and another of size $q+k$. The rate coefficient for this reaction is $a(p,q;k)\geqslant 0.$
Assuming the validity of the mass action law of chemical kinetics, the rate of production of $(p-k)$-clusters and $(q+k)$-clusters due to the reaction between $p$-clusters and $q$-clusters, according to the mechanism displayed in Figure 1, is equal to $a(p,q;k)c_{p}(t)c_{q}(t),$ where the rate coefficient is $a(p,q;k).$
@@
In scheme (1), one $i$-cluster is created if (i) $p-k=i$, or if (ii) $q+k=i.$ Likewise, one $i$-cluster is destroyed if (iii) $q=i$, or if (iv) $p=i.$ In the case of creation (i), for every $k\in{\mathbb{N}}_{+}:={\mathbb{N}}\setminus\{0\},$ we have to consider all the indices combinations for which $p=i+k$ and $q=j,$ for any $j\in{\mathbb{N}}.$ Therefore, in this case, the formation of an $i$-cluster proceeds through the scheme (1) illustrated in Figure 1 and, by the mass action law, its contribution to the global rate of change of $c_{i}(t)$ is given by
@@
For the creation case (ii), we have to consider all possible indices combinations for which $q=i-k,$ that is, for $1\leqslant k\leqslant i,$ we have $p\geqslant k,$ which, by calling $j:=p$ leads to the contribution
@@
If the system is not isolated but instead the number density of 00-clusters is kept constant at some fixed value $c_{0}(t)=c_{00}$ the equation for the evolution of $c_{0}(t)$ becomes simply
@@
Condition (10) is due to the fact that, as pointed out above, each rate coefficient $a(i,j;k)$ with $i=k$ corresponds to the rate of a coagulation reaction of a $k$-cluster with a $j$-cluster (to produce a $(j+k)$-cluster) and thus this is the usual symmetry assumption due to the fact that a coagulation reaction of a $k$ with a $j$-cluster is the same as the reaction of a $j$ with a $k$-cluster. Similarly, (11) correspond to the symmetry of the rates of the reaction $\langle i\rangle+\langle 0\rangle\rightarrow\langle i-k\rangle+\langle k\rangle$, when interpreted as a fragmentation of the $i$-cluster produced by the shedding off a $k$-cluster, or by the shedding off a $(i-k)$-cluster.
@@
As was already pointed out in section 2, it is interesting to observe that, if the rate coefficients satisfy
@@
and the symmetry conditions (10)-(11) hold, then DGED system in the non-isolated case is reduced to the usual discrete coagulation-fragmentation equations. In fact, observe that with these conditions we get
$Q_{1,i}(c)$ ==(14)${=\joinrel=}\sum_{k=1}^{\infty}a(i+k,0;k)c_{i+k}c_{0}$ (15)
$Q_{2,i}(c)$ ==(13)${=\joinrel=}-\sum_{k=1}^{\infty}a(k,i;k)c_{k}c_{i}$ (16)
$Q_{3,i}(c)$ ====(13),(14)${=\joinrel=\joinrel=\joinrel=}\sum_{j=i}^{\infty}a(j,0;i)c_{j}c_{0}+\sum_{k=1}^{i}a(k,i-k;k)c_{k}c_{i-k}$ (17)
$Q_{4,i}(c)$ ====(13),(14)${=\joinrel=\joinrel=\joinrel=}-\sum_{k=1}^{i}a(i,0;k)c_{i}c_{0}-\sum_{j=0}^{\infty}a(i,j;i)c_{i}c_{j}$ (18)
@@
$\sum_{j=1}^{\infty}a(i,j;i)c_{i}c_{j}$ ==(10)${=\joinrel=}\sum_{j=1}^{\infty}a(j,i;j)c_{i}c_{j}$ (19)
$\sum_{j=i+1}^{\infty}a(j,0;i)c_{j}c_{0}$ ==(11)${=\joinrel=}\sum_{j=i+1}^{\infty}a(j,0;j-i)c_{j}c_{0}=\sum_{k=1}^{\infty}a(i+k,0;k)c_{i+k}c_{0},$ (20)
@@
where $W_{i,j}(c):=2a(i,j;i)c_{i}c_{j}-2a(i+j,0;j)c_{i+j}c_{0}$. Clearly, if our system is immersed in an infinite particle bath of 00-cluster particles (which means that the concentration of 00-particles is kept constant) and if we define $a_{i,j}:=2a(i,j;i)$ and $b_{i,j}:=2a(i+j,0;j)c_{0}$, then (21) becomes the usual discrete coagulation-fragmentation equation [8]. Therefore, in this case we have
@@
3. Mathematical setting
@@
$X_{0,1}:=\bigl{\{}c:{\mathbb{N}}\to{\mathbb{R}}\,|\;\|c\|:=\|c\|_{\ell_{1}}+\|(ic_{i})\|_{\ell_{1}}<\infty\bigr{\}},$ (25)
where $\|u\|_{\ell_{1}}:=\sum_{i=0}^{\infty}|u_{i}|$ is the usual $\ell_{1}$ norm of the sequence $u:{\mathbb{N}}\to{\mathbb{R}}$.
@@
We are now ready to state the following
@@
With respect to the bounds on the rate coefficients, we assume the existence of positive constants $C$ and $\mathcal{Q}$, which as a matter of convenience we take $C,\mathcal{Q}\geqslant 1$, as well as non-negative numbers $q_{i,k},$ $1\leqslant k\leqslant i,$ verifying
@@
thus resulting in ${\displaystyle\sum_{k=1}^{i}q_{i,k}\leqslant\mathcal{Q}}$, and, in particular,
@@
Remark that for the exchange driven growth system with kernel verifying $K(i,j)\leqslant C_{0}ij$, for $i,j\in{\mathbb{N}}_{+},$ by (12), we have, for all $i,j\in{\mathbb{N}}_{+}$
@@
which is (28) with $C=C_{0}$ and $q_{i,k}=\delta_{k,1},$ observing that (27) is also verified since,
@@
In fact, this is a particular case of the more general one for which there is an upper bound, $\bar{k}\geqslant 1,$ for the number of particles exchanged between two reacting clusters of any sizes, that is,
@@
An example of a strictly positive rate coefficient $a(i,j;k)$ without a fixed upper bound $\bar{k}$ on the size of clusters being exchanged is the constant kernel $a(i,j;k)=1$, for which we have
@@
A less trivial example of the same sort but such that the rate coefficients are unbounded is
@@
Finally, for the coagulation-fragmentation case with upper bound on the coagulation rate coefficients of the type $a_{i,j}\leqslant C_{1}(i+j),$ by (22), we have,
@@
We therefore conclude that the cases of the exchange driven system growth with kernel of multiplicative bound type and the coagulation-fragmentation system with coagulation kernel with linear additive bound are both particular cases of the discrete generalized exchange driven system with bounds of the type (27)-(28).
Prior to establish the existence result in section 5, it is imperative to revisit certain notations and to get some inequalities that will be needed later. Let us first recall the space $\mathcal{E}$ from [18], defined as the set of non-negative and convex functions $\sigma\in\mathcal{C}^{1}([0,\infty))\cap W^{2,\infty}_{loc}(0,\infty)$ with $\sigma(0)=0$, $\sigma^{\prime}(0)\geqslant 0$ and $\sigma^{\prime}$ is a concave function. Assume furthermore that these functions satisfy the following condition:
@@
Next, we define $\mathcal{E}_{1}$ to be the set of all non-negative and convex functions $\sigma\in\mathcal{C}^{2}([0,\infty))$ with $\sigma(0)=0$, $\sigma^{\prime}(0)=0$, and $\sigma^{\prime}$ is a convex function satisfying the $\Delta_{2}-$condition, that is, there is a constant $A_{\sigma}\geqslant 0,$ such that
@@
For example: if $\sigma(x)=x^{p}$, then $\sigma\in\mathcal{E}$ if $p\in(1,2]$, and $\sigma\in\mathcal{E}_{1}$ if $p\geqslant 2$.
@@ 4. The truncated system
[…]
@@
The following two lemmas about properties of the solutions to the truncated system, in which we assume the conditions (27)-(28), will be essential for the proof of the existence theorem in the next section. They use, in a crucial way, the inequalities proved in lemmas 3.2 and 3.3.
@@
[…]
@@
In order to prove (52), we integrate both members of (61) and take in account (51) and the sign of each term, thus getting
[…]
@@
[…]
Since $-\tilde{\sigma}(i,0,k)=|\tilde{\sigma}(i,0,k)|=\sigma(i)-\sigma(i-k)-\sigma(k),$ the conclusion is obtained from (4) and (64) by application of Lemma 3.3 and redefining the constant $\gamma_{T}.$

### v2

@@ Abstract.
Corresponding author: P. K. Barik
@@ 1. Introduction
In the present paper our primary goal is to propose a new, more general, discrete model for exchange-driven growth of cluster and start its mathematical study. This new model will be referred to as the discrete generalized exchange-driven growth model (DGED, for short). As explained in the next session, the DGED model allows for the exchange not only of a monomer between two interacting clusters, but of the transfer of a bigger chunk of one cluster to the other. Thus, the DGED model includes as a special case the EDG model referred to above. In a certain sense, the usual coagulation-fragmentation equations (with binary fragmentation) can also be considered as a special case of the DGED equations.
@@
The outline of this article is as follows. In sections 2 and 3, we introduce the DGED model for two cases that, in section 2, are called isolated and non-isolated, and some relevant mathematical notions and results that will be needed later in this work. In addition, all the main results of our paper are stated in section 3. In section 4 we consider a truncated finite dimensional version of the DGED system. This is a finite dimensional ordinary differential equation system, for which the existence, uniqueness, and non-negativity of solutions to Cauchy problems are automatically obtained from the standard theory. Here we prove some properties of its solutions that will play an important role in the next section. In section 5 we establish the existence of mild solutions to Cauchy problems for the DGED system, i.e., continuous solutions of the integral version of the differential equations system. This is done by the use of Helly’s theorem and a diagonal argument to establish the existence of a function which is the limit of a (sub)sequence of solutions to truncated systems as the dimension of the truncation grows to infinity, and then by proving that this function is indeed a (mild) solution of the DGED system using appropriate bounds (obtained in the previous two sections) and the bounded and monotone convergence theorems. Still in section 5 we prove that for the isolated case solutions obtained in this way (which are called admissible) conserve two important quantities, which have the “physical” interpretation of being the total number of clusters, and the total mass of the system, whereas in the non-isolated case only mass is conserved. In the short section 6 we prove that under slightly more restrictive conditions on the rate coefficients of the DGED system and assuming conservation of the total cluster mass (which is true for admissible solutions in both the isolated and non-isolated cases), then each component of the mild solution proved to exist in section 5 is indeed continuously differentiable. In Section 7, we prove a uniqueness result under a more restrictive growth condition on the rate coefficients, applying standard techniques used in coagulation-type systems. We conclude the paper with a section of final remarks where we discuss some possible future work related to the long-term behavior of solutions for this model.
@@ 2. The generalized exchange-driven growth model
Figure 1. Reaction scheme of the DGED model considered in this paper. A chunk of size $k$ in a particle of size $p$ is transferred to a particle of size $q$ to produce a particle of size pkp-k and another of size $q+k$. The rate coefficient for this reaction is $a(p,q;k)\geqslant 0.$
Assuming the validity of the mass action law of chemical kinetics, the rate of production of (pk)(p-k)-clusters and $(q+k)$-clusters due to the reaction between $p$-clusters and $q$-clusters, according to the mechanism displayed in Figure 1, is equal to $a(p,q;k)c_{p}(t)c_{q}(t),$ where the rate coefficient is $a(p,q;k).$
@@
In scheme (1), one $i$-cluster is created if (i) pk=ip-k=i, or if (ii) $q+k=i.$ Likewise, one $i$-cluster is destroyed if (iii) $q=i$, or if (iv) $p=i.$ In the case of creation (i), for every $k\in{\mathbb{N}}_{+}:={\mathbb{N}}\setminus\{0\},$ we have to consider all the indices combinations for which $p=i+k$ and $q=j,$ for any $j\in{\mathbb{N}}.$ Therefore, in this case, the formation of an $i$-cluster proceeds through the scheme (1) illustrated in Figure 1 and, by the mass action law, its contribution to the global rate of change of $c_{i}(t)$ is given by
@@
For the creation case (ii), we have to consider all possible indices combinations for which q=ik,q=i-k, that is, for $1\leqslant k\leqslant i,$ we have $p\geqslant k,$ which, by calling $j:=p$ leads to the contribution
@@
If the system is not isolated but instead the number density of 0-clusters is kept constant at some fixed value $c_{0}(t)=c_{00}$ the equation for the evolution of $c_{0}(t)$ becomes simply
@@
Condition (10) is due to the fact that, as pointed out above, each rate coefficient $a(i,j;k)$ with $i=k$ corresponds to the rate of a coagulation reaction of a $k$-cluster with a $j$-cluster (to produce a $(j+k)$-cluster) and thus this is the usual symmetry assumption due to the fact that a coagulation reaction of a $k$ with a $j$-cluster is the same as the reaction of a $j$ with a $k$-cluster. Similarly, (11) correspond to the symmetry of the rates of the reaction $\langle i\rangle+\langle 0\rangle\rightarrow\langle i-k\rangle+\langle k\rangle$, when interpreted as a fragmentation of the $i$-cluster produced by the shedding off a $k$-cluster, or by the shedding off a (ik)(i-k)-cluster.
@@
As was already pointed out above, it is interesting to observe that the DGED system formally reduces to the standard discrete coagulation-fragmentation equations in the non-isolated case, provided the rate coefficients satisfy the conditions
@@
In fact, observe that with these conditions we get
[…]
@@
[…]
@@
where $W_{i,j}(c):=2a(i,j;i)c_{i}c_{j}-2a(i+j,0;j)c_{i+j}c_{0}$. Clearly, if our system is immersed in an infinite particle bath of 0-cluster particles (which means that the concentration of 0-particles is kept constant) and if we define $a_{i,j}:=2a(i,j;i)$ and $b_{i,j}:=2a(i+j,0;j)c_{0}$, then (21) becomes the usual discrete coagulation-fragmentation equation [8]. Therefore, in this case we have
@@
3. Mathematical setting and main results
@@
$X_{0,1}:=\Bigl{\{}c:{\mathbb{N}}\to{\mathbb{R}}\,\bigl{|}\;\|c\|:=\sum_{i=0}^{\infty}(1+i)|c_{i}|<\infty\Bigr{\}}.$ (25)
Note that we can write $\|c\|=\|c\|_{\ell_{1}}+\|(ic_{i})\|_{\ell_{1}}$ where $\|u\|_{\ell_{1}}$ is the usual $\ell_{1}$ norm of the sequence $u=(u_{i}):{\mathbb{N}}\to{\mathbb{R}}$.
@@
We are now ready to state the following definition of solution.
@@
3.1. Rate Coefficients: Growth Assumptions and Examples
As stated previously, we assume conditions (9), (10) and (11) throughout the paper. With respect to the bounds on the rate coefficients, we assume the existence of positive constants $C$ and $\mathcal{Q}$, which as a matter of convenience we take $C,\mathcal{Q}\geqslant 1$, as well as non-negative numbers $q_{i,k},$ with $1\leqslant k\leqslant i,$ satisfying
@@ 3.1. Rate Coefficients: Growth Assumptions and Examples
We can see $q_{i,k}$ as measuring the ease with which a cluster of size $k$ can be detached from a cluster of size $i$: the smaller the value of $q_{i,k}$ is, the smaller is the rate coefficient $a(i,j;k)$, and so the harder it is for a $k$-cluster to take part in the exchange reaction $\langle i\rangle+\langle j\rangle\to\langle i-k\rangle+\langle j+k\rangle.$
@@
thus resulting in $\sum_{k=1}^{i}q_{i,k}\leqslant\mathcal{Q}$}, and, in particular,
@@
We illustrate next by exhibiting some examples that (27) and (28) are fulfilled by some well known cases in the literature.
Example 1. Consider the exchange driven growth system with kernel satisfying $K(i,j)\leqslant C_{0}ij$, for $i,j\in{\mathbb{N}}_{+},$ by (12), we have, for all $i,j\in{\mathbb{N}}_{+}$
@@
which is (28) with $C=C_{0}$ and $q_{i,k}=\delta_{k,1},$ observing that (27) is also satisfied since,
@@
Example 2. In fact, Example 1 is a particular case of the more general case where there is an upper bound, $\bar{k}\geqslant 1,$ for the number of particles exchanged between two reacting clusters of any sizes, that is,
@@
Example 3. An example of a strictly positive rate coefficient $a(i,j;k)$ without a fixed upper bound $\bar{k}$ on the size of clusters being exchanged is the constant kernel $a(i,j;k)=1$, for which we have
@@
Example 4. A less trivial example of the sort considered in Example 3 but such that the rate coefficients are unbounded is
@@
Example 5. Finally, in this example let us consider the coagulation-fragmentation case with upper bound on the coagulation rate coefficients of the type $a_{i,j}\leqslant C_{1}(i+j),$ by (22). For this case we have
@@
Therefore, from these examples (in particular from examples 1 and 5) we conclude that the cases of the exchange driven system growth with kernel of multiplicative bound type and the coagulation-fragmentation system with coagulation kernel with linear additive bound are both particular cases of the discrete generalized exchange driven system with bounds of the type (27)-(28).
3.2. Main results
In this subsection, we state the main results proved in this paper for the initial value problem (7)–(8), considering both the isolated and non-isolated cases.
Let us first state the existence result for the isolated case.
Theorem 3.2.
Let $c_{0}=(c_{0i})_{i\geq 0}\in X_{0,1}^{+}$. Assume the rate coefficients satisfy the conditions (27)–(28). Then the initial value problem (7), (8) for the isolated case has at least one global solution $c$ defined on the interval $0,\infty)$.
The next theorem establishes mass and particle number conservation in the isolated system.
Theorem 3.3.
and
Similarly to Theorem 3.2, we now state the existence of global solutions for the non-isolated case.
Theorem 3.4.
Let $c_{0}=(c_{0i})_{i\geq 0}\in X_{0,1}^{+}$. Assume the rate coefficients satisfy the conditions (27)–(28). Then the initial value problem (7), (8) for the non-isolated case has at least one global solution $c$ defined on the interval $0,\infty)$.
The following result establishes the mass conservation property in the non-isolated case:
Theorem 3.5.
We next state the regularity of the solution in classical sense:
Theorem 3.6 (Regularity of solutions).
[…]
We finally state a partial uniqueness result:
Theorem 3.7.
3.3. Preliminary results
Prior to establish the existence result in section 5, it is imperative to revisit certain notations and to get some inequalities that will be needed later. Let us first recall the space $\mathcal{E}$ from [18]:
$\mathcal{E}$ is the set of non-negative and convex functions $\sigma\in\mathcal{C}^{1}([0,\infty))\cap W^{2,\infty}_{loc}(0,\infty)$ with $\sigma(0)=0$, $\sigma^{\prime}(0)\geqslant 0$ and $\sigma^{\prime}$ is a concave function. Assume furthermore that these functions satisfy the following condition:
@@ 3.3. Preliminary results
Again from [18], define $\mathcal{E}_{1}$ as follows:
$\mathcal{E}_{1}$ is the set of all non-negative and convex functions $\sigma\in\mathcal{C}^{2}([0,\infty))$ with $\sigma(0)=0$, $\sigma^{\prime}(0)=0$, and $\sigma^{\prime}$ is a convex function satisfying the so-called $\Delta_{2}$-condition, namely, there is a constant $A_{\sigma}\geqslant 0,$ such that
@@
As an illustration of these concepts consider the following example: if $\sigma(x)=x^{p}$, then $\sigma\in\mathcal{E}$ when $p\in(1,2]$, and $\sigma\in\mathcal{E}_{1}$ if $p\geqslant 2$.
@@ 4. The truncated system
$\left\{\begin{aligned} \dot{c}_{0}&=\sum_{j=1}^{2}Q_{j,0}^{N}(c)\\ \dot{c}_{i}&=\sum_{j=1}^{4}Q_{j,i}^{N}(c),\quad i\in\{1,\ldots,N-1\}\\ \dot{c}_{N}&=\sum_{j=3}^{4}Q_{j,N}^{N}(c),\end{aligned}\right$. (38)
@@
The following two lemmas about properties of the solutions to the truncated system, in which we assume the conditions (27)-(28), will be essential for the proof of the existence theorem in the next section. They use, in a crucial way, the inequalities proved in lemmas 3.8 and 3.9.
@@
$\mathcal{J}^{N}_{0}:=\{(p,k)\in({\mathbb{N}}_{+})^{2}:M_{0}\leqslant p\leqslant N,\;1\leqslant k\leqslant p-1\},$
@@
In order to prove (54), we integrate both members of (63) and take into account (53) and the sign of each term, thus getting
$\int_{0}^{T}\Big{|}\sum_{k=1}^{N-1}\sum_{i=k}^{N}\tilde{\sigma}(i,0,k)a(i,0;k)c_{0}^{N}(t)c_{i}^{N}(t)\Big{|}dt\leqslant$ $\sum_{i=0}^{N}\sigma(i)c^{N}_{i}(0)+K\int_{0}^{T}\sum_{i=0}^{N}\sigma(i)c_{i}^{N}(t)dt$
@@
$\int_{0}^{T}\Big{|}\sum_{k=1}^{N-1}\sum_{i=k}^{N}\tilde{\sigma}(i,0,k)a(i,0;k)c_{0}^{N}(t)c_{i}^{N}(t)\Big{|}dt\\ =\int_{0}^{T}\Big{|}\sum_{i=1}^{N}\sum_{k=1}^{i-1}\tilde{\sigma}(i,0,k)a(i,0;k)c_{0}^{N}(t)c_{i}^{N}(t)\Big{|}dt.$ (66)
Since $\tilde{\sigma}(i,0,k)=|\tilde{\sigma}(i,0,k)|=\sigma(i)-\sigma(i-k)-\sigma(k),$ the conclusion is obtained from (4) and (66) by application of Lemma 3.9 and redefining the constant $\gamma_{T}.$